OpenStudy (anonymous):

A point C is chosen on the line segment AB such that AC/BC =(BC)/(5AB). Find (AC)/(BC)

OpenStudy (anonymous):

no that was a mistake, let me start again

OpenStudy (anonymous):

ok that looks better. i think it is \[\frac{1}{\sqrt{5}}\]

OpenStudy (anonymous):

 i made AB = 1, AC = x and so CB = 1-x then i solved \[\frac{x}{1-x}=\frac{1-x}{5x}\] \[5x^2=(1-x)^2\] \[x=\frac{\sqrt{5}-1}{4}\] then took the ratio \[\frac{\frac{\sqrt{5}}{4}}{1-\frac{-1+\sqrt{5}}{4}}\]

OpenStudy (anonymous):

\[\frac{\sqrt{5}}{4-(-1+\sqrt{5})}\] \[\frac{-1+ \sqrt{5}}{5-\sqrt{5}}\] \[\frac{1}{\sqrt{5}}\]

OpenStudy (anonymous):

typo there, first line should be \[\frac{-1+\sqrt{5}}{4-(-1+\sqrt{5})}\]

OpenStudy (anonymous):

\[5x ^{2}=(1-x)^{2}\] \[5x ^{2}=1-2x+x^{2}\] \[4x ^{2}+2x-1=0\] By simplifying: \[x=(-1+\sqrt{2})\div4\] For AC/BC \[((-1+\sqrt{2})\div4)\div(1-((-1+\sqrt{2})\div4))\] \[=(-7+6\sqrt{2})\div3\]

OpenStudy (anonymous):

i still think it is \[5x^2=(1-x)^2\] \[x=\frac{-1+\sqrt{5}}{4}\] but i may be wrong

OpenStudy (anonymous):

no, i am not wrong, it is right http://www.wolframalpha.com/input/?i=4x^2%2B2x-1%3D0

OpenStudy (anonymous):

Yeah you are right. Forgot to multiply the "a" part in the formula b^2-4ac. That is just one silly mistake. Oh boy. :P

OpenStudy (phi):

There is something I don't like about this problem AC/BC =(BC)/(5AB) with Sat's definitions: AB = 1, AC = x and so CB = 1-x isn't the ratio: \[ \frac{x}{1-x}= \frac{1-x}{5} \] Note there is no x in the denominator on the right hand side. But this ratio leads to negative numbers for x.

OpenStudy (anonymous):

@phi Now that you mention it. Your formula is right. There is no 5x involved since 5AB=5.

OpenStudy (phi):

As an exercise I finished it off this way: 5x= 1-2x+x^2 or x^2-7x+1= 0 this has two roots: (7± 3sqrt(5))/2 of which one leads to positive x: x= (7-3sqrt(5))/2 and 1-x = (3sqrt(5)-5)/2 the ratio of x/(1-x) = (3sqrt(5)-5)/10 = 0.17082